OpenStudy (anonymous):

A 57.0 kg diver dives from a height of 15.0 m. She reaches a speed of 14.0 m/s just before entering the water. a) What was the average force of air resistance (e.g., friction) acting on the diver? b) What is the force of friction underwater if she reaches a depth of 2.5 m before stopping? Don't neglect the buoyant force of 500 N acting on the diver once underwater

OpenStudy (anonymous):

I know part a. Can someone help me in b please?

OpenStudy (anonymous):

a) \[F_r=m*a\] \[F_r=W-F_{res}\] \[V_f^2=V_o^2+2*a*\Delta S\] \[a=6.53~m/s^2\] \[F_{res}=m*g-m*a\] \[F_{res}=m*(g-a)\] \[F_{res}=57*(9.8-6.53)\] \[\boxed{\boxed{F_{res}=186.39~N}}\] b) \[F_r=m*a\}\] \[F_r=W-E-F_{res}\] \[F_{res}=W-E-F_{r}\] \[V_o^2=2*a*\Delta S\] \[a=39.2~m/s^2\] \[F_{res}=m*g-E-m*a\] \[\boxed{\boxed{F_{res}=-2175.8~N}}\]